The cost of three notebooks and four pencils is $8.50. The cost of five notebooks and eight pencils is $14.50. Determine the cost of one notebook and the cost of one pencil.
Answer by Cromlix(4381)

Notebook = x
Pencil = y
3x + 4y = 8.50........1
5x + 8y = 14.50.......2
Multiply (1) by 2
6x + 8y = 17.00.......3
5x + 8y = 14.50.......2
Subtract (2) from (3)
x = 2.50
Substitute x = 2.50 in Eq(1)
3x + 4y = 8.50
7.50 + 4y = 8.50
4y = 8.50 - 7.50
4y = 1.00
y = 0.25
Notebook costs $2.50
Pencil costs $0.25
Hope this helps.
:-)